# CDsampling: An R Package for Constrained D-Optimal Sampling in Paid Research Studies

### v3

@@ 4   Summary
where $\mu_{i}=E(Y_{i}\mid{\mathbf{x}}_{i})$, $Y_{i}\in\{0,1\}$, ${\mathbf{x}}_{i}=(x_{i1},x_{i2})^{\top}\in\{(-1,-1),(-1,1),(1,-1)\}$ and parameters $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\beta_{2})=(0.5,0.5,0.5)$. The approximate allocation to the three possible design points, namely, ${\mathbf{x}}_{i}$’s, is $\mathbf{w}=(1/3,1/3,1/3)^{\top}$. To calculate this Fisher information matrix $\mathbf{F}$ given the design ${\mathbf{w}}$, we may use F_func_GLM() function in the CDsampling package. Additionally, W_func_GLM() function can be used to find the diagonal elements of the matrix $\mathbf{W}$ in the Fisher information matrix (1) of GLM.
@@
We revisit the trauma_data example previously analyzed in Section 3.2. Consider a research study under a cumulative logit npo model with $J=5$ response levels and covariates $(x_{i1},x_{i2})\in\{(1,0),(2,0),(3,0),(4,0),(1,1),(2,1),(3,1),$ (4,1)}(4,1)\}( 4 , 1 ) }. The model can be written as the following four equations:
@@
where $i=1,\dots,8$. We assume that the true parameters $\boldsymbol{\beta}=(\hat{\beta}_{11},\hat{\beta}_{12},\hat{\beta}_{13},\hat{\beta}_{21},\hat{\beta}_{22},\hat{\beta}_{23},\hat{\beta}_{31},$ $\hat{\beta}_{32},\hat{\beta}_{33},\hat{\beta}_{41},\hat{\beta}_{42},\hat{\beta}_{43})^{\top}$ $=$ (4.047,0.131,4.214,2.225,0.376,3.519,0.302,0.237,2.420,1.386,(-4.047,-0.131,4.214,-2.225,-0.376,3.519,-0.302,-0.237,2.420,1.386,( - 4.047 , - 0.131 , 4.214 , - 2.225 , - 0.376 , 3.519 , - 0.302 , - 0.237 , 2.420 , 1.386 , $-0.120,1.284)^{\top}$, and the approximate allocation for the eight design points is $\mathbf{w}=(1/8,\ldots,1/8)^{\top}$.
@@
> J=5; p=12; m=8;> beta = c(-4.047, -0.131, 4.214, -2.225, -0.376, 3.519, -0.302, -0.237,  2.420, 1.386,-0.120,  1.284)> Xi=rep(0,J*p*m); dim(Xi)=c(J,p,m)> Xi[,,1] = rbind(c( 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 1, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 1, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,2] = rbind(c( 1, 2, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 2, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 2, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 2, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,3] = rbind(c( 1, 3, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 3, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 3, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 3, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,4] = rbind(c( 1, 4, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 4, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 4, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 4, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,5] = rbind(c( 1, 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 1, 1, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 1, 1, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 1),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,6] = rbind(c( 1, 2, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 2, 1, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 2, 1, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 2, 1),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,7] = rbind(c( 1, 3, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 3, 1, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 3, 1, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 3, 1),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,8] = rbind(c( 1, 4, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 4, 1, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 4, 1, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 4, 1),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> alloc = rep(1/8,m)> F_func_MLM(w=rep(1/8,m), beta=beta, X=Xi, link=’cumulative’)Dimensions: 12 x 12Matrix:------------------------------------------------------------      [,1]        [,2]        [,3]        [,4]        [,5]        [,6] [1,]  0.44505694  1.37915564  0.43609135 -0.37247296 -1.21252053 -0.36379349 [2,]  1.37915564  4.78410934  1.35694145 -1.21252053 -$4$1.19085831 [3,]  0.43609135  1.35694145  0.43609135 -0.36379349 -1.19085831 -0.36379349 [4,] -0.37247296 -1.21252053 -0.36379349  0.51192600  1.55177413  0.48018678 [5,] -1.21252053 -$4$1.19085831  1.55177413  5.31193908  1.48241981 [6,] -0.36379349 -1.19085831 -0.36379349  0.48018678  1.48241981  0.48018678 [7,]  0.00000000  0.00000000  0.00000000 -0.09154268 -0.22027625 -0.07471168 [8,]  0.00000000  0.00000000  0.00000000 -0.22027625 -0.64323991 -0.18445802 [9,]  0.00000000  0.00000000  0.00000000 -0.07471168 -0.18445802 -0.07471168[10,]  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000[11,]  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000[12,]  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000      [,7]        [,8]        [,9]        [,10]       [,11]       [,12] [1,]  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000 [2,]  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000 [3,]  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000 [4,] -0.09154268 -0.22027625 -0.07471168  0.00000000  0.00000000  0.00000000 [5,] -0.22027625 -0.64323991 -0.18445802  0.00000000  0.00000000  0.00000000 [6,] -0.07471168 -0.18445802 -0.07471168  0.00000000  0.00000000  0.00000000 [7,]  0.29320484  0.71978889  0.16205254 -0.10435894 -0.25399393 -0.06194131 [8,]  0.71978889  2.13312603  0.41294396 -0.25399393 -0.74842743 -0.15305771 [9,]  0.16205254  0.41294396  0.16205254 -0.06194131 -0.15305771 -0.06194131[10,] -0.10435894 -0.25399393 -0.06194131  0.17861575  0.45287619  0.06925715[11,] -0.25399393 -0.74842743 -0.15305771  0.45287619  1.37180187  0.17395849[12,] -0.06194131 -0.15305771 -0.06194131  0.06925715  0.17395849  0.06925715------------------------------------------------------------
@@
This suggests that in the feasible allocation set $S_{0}=\{(w_{1},w_{2},w_{3})^{\top}$in $\mathbb{R}^{3}\mid w_{i}\geq 0,i=1,\ldots,3;\sum_{i=1}^{3}w_{i}=1\}$, the D-optimal allocation is $\mathbf{w}=(w_{1},w_{2},w_{3})=(1/3,1/3,1/3)$. $\Box$
@@
We also need to find the upper boundary and lower boundary functions for $[r_{i1},r_{i2}]$ in step $3^{\circ}$ of the constrained lift-one algorithm to meet the conditions in $S$. The $r_{i1}$ and $r_{i2}$ can be obtained as follows:
@@
To determine the EW D-optimal allocation with the independent uniform priors $\beta_{0}\sim$ uniform(2,2)22(-2,2)( - 2 , 2 ), $\beta_{1}\sim$ uniform$(-1,5)$, $\beta_{21}\sim$ uniform$(-1,5)$, and $\beta_{22}\sim$ uniform(-1,5), as outlined in Section 3.1, we utilize the same codes provided in Section 3.1:
> library(cubature)> unif.prior <- rbind(c(-2, -1, -1, -1), c(2,  5,  5, 5))> W.EW.unif = matrix(rep(0,6))> for (i in 1:6){+   x = matrix((cbind(1, unique(X)))[i,])+   W.EW.unif[i] = hcubature(function(beta) dunif(beta[1], min=unif.prior[1,1],max=unif.prior[2,1])*dunif(beta[2], min=unif.prior[1,2],max=unif.prior[2,2])*dunif(beta[3], min=unif.prior[1,3],max=unif.prior[2,3])*dunif(beta[4], min=unif.prior[1,4], max=unif.prior[2,4])*(exp(x[1]*beta[1]+x[2]*beta[2]+x[3]*beta[3]+x[4]*beta[4])/(1+exp(x[1]*beta[1]+x[2]*beta[2]+x[3]*beta[3]+x[4]*beta[4]))^2), lowerLimit = unif.prior[1,],upperLimit  = unif.prior[2,])\$integral+ }> set.seed(602)> approximate_design_EW = liftone_constrained_GLM(X=X, W=W.EW.unif,  g.con=g.con,+ g.dir=g.dir, g.rhs=g.rhs, lower.bound=lower.bound, upper.bound=upper.bound,+ reltol=1e-12, maxit=100, random=TRUE, nram=12, w00=NULL, epsilon=1e-12)> exact_design_EW = approxtoexact_constrained_func(n=200, w=approximate_design_EW\$w,+ m=6, beta=beta, link=’logit’, X=X, Fdet_func=Fdet_func_GLM, iset_func=iset_func_trial)
@@
> nsimu=100> p=3 # 3 covariates (excluding intercept)> set.seed(666)> seeds <- sample(1:10000, nsimu)   #random seeds> # matrix for recording beta estimates> beta.estimate <- rep(NA, nsimu*(p+1)*(5))> dim(beta.estimate) = c(nsimu, p+1, 5) # [1] 100   4   8> dimnames(beta.estimate)[[2]] = paste0("beta",0:p,seq="")> dimnames(beta.estimate)[[3]] <- c(’full’, "SRSWOR", "Unif", "local_Dopt", "EW_Unif")> # generate X for original 500 patients> # gender group (0 for female and 1 for male)> gender = c(rep(0,50),rep(0,40),rep(0,10),rep(1,200),rep(1,150),rep(1,50))> # age group (0 for 18~25, 1 for 26~64, and 2 for 65 or above)> age1 = c(rep(0,50), rep(1,40), rep(0,10), rep(0,200), rep(1,150), rep(0,50))> age2 = c(rep(0,50), rep(0,40), rep(1,10), rep(0,200), rep(0,150), rep(1,50))> X.original = cbind(gender, age1, age2)> #probability of Y = 1 given X using coefficient beta under logistic model> prob = exp(beta %*t(cbind(1,X.original))) / (1+exp(beta %*t(cbind(1,X.original))))> label = rep(0, length(X.original[,1])) #create label to store category of the population data> for(k in 1:length(X.original[,1])){+   for (m in 1:m){+     if((X.original[k,]==unique(X.original)[m,])[1] & (X.original[k,]==unique(X.original)[m,])[2]& (X.original[k,]==unique(X.original)[m,])[3]){+       label[k] = m+     }+   }+ }> n = tabulate(label) #number of subjects available in each category> sampling_func = function(label, n.sample, X, s, n, m){+   J = 0+   j = rep(0, m)+   c = rep(0, m)+   isample = vector()+   while(J < n.sample){+     for(k in 1:length(X[,1])){+       i = label[k]+       c[i] = c[i] + 1+       if(j[i]<s[i]){+         if(runif(1) < (s[i] - j[i])/(n[i] - c[i] + 1)){+           j[i]=j[i]+1+           J = J + 1+           isample = c(isample, k)+         }+       }+     }+   }+   return(isample)+ }> s_unif = unif_design\$allocation #num to sample in constrained uniform from each category> s_localD = exact_design\$allocation #num to sample in local D-optimal from each category> s_EWD = exact_design_EW\$allocation #num to sample in EW D-optimal from each category> for(isimu in 1:nsimu){+   set.seed(seeds[isimu])+   ##Generate Y response under logistic model+   Y = rbinom(n=length(X.original[,1]), size=1, prob = prob)+   #Estimate beta with full data+   beta.estimate[isimu,,"full"] <- glm(Y~X.original,family = "binomial")\$coefficients+   #SRSWOR sample+   isample1 <- sample(x=1:500, size=200); #simple sample without replacement, output selected label+   X1=X.original[isample1,];+   Y1=Y[isample1];+   #Estimate beta with SRSWOR sampling+   beta.estimate[isimu,,"SRSWOR"] <- glm(Y1~X1,family = "binomial")\$coefficients+   #Uniform sample+   #stratified function returns s[k] (num of samples plan to collect for ith category)+   isample2 = sampling_func(label=label, n.sample=200, X=X.original, s=s_unif, n=n, m=m)+   X2=X.original[isample2,];+   Y2=Y[isample2];+   #Estimate beta with Unif sampling+   beta.estimate[isimu,,"Unif"] <- glm(Y2~X2,family = "binomial")\$coefficients+   #Local D-opt Sample+   isample3 = sampling_func(label=label, n.sample=200, X=X.original, s=s_localD, n=n,m=m)+   X3=X.original[isample3,];+   Y3=Y[isample3];+   #Estimate beta with local D-optimal sampling+   beta.estimate[isimu,,"local_Dopt"] <- glm(Y3~X3,family = "binomial")\$coefficients+   #EW D-opt Sample+   isample4 = sampling_func(label=label, n.sample=200, X=X.original, s=s_EWD, n=n, m=m)+   X4=X.original[isample4,];+   Y4=Y[isample4];+   #Estimate beta with EW D-optimal sampling+   beta.estimate[isimu,,"EW_Unif"] <- glm(Y4~X4,family = "binomial")\$coefficients+ }

### v4

@@ 4  Summary
where $\mu_{i}=E(Y_{i}\mid{\mathbf{x}}_{i})$, $Y_{i}\in\{0,1\}$, ${\mathbf{x}}_{i}=(x_{i1},x_{i2})^{\top}\in\{(-1,-1),(-1,1),(1,-1)\}$ and parameters $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\beta_{2})^{\top}=(0.5,0.5,0.5)^{\top}$. The approximate allocation to the three possible design points, namely, ${\mathbf{x}}_{i}$’s, is $\mathbf{w}=(1/3,1/3,1/3)^{\top}$. To calculate this Fisher information matrix $\mathbf{F}$ given the design $\mathbf{w}$}, we may use F_func_GLM() function in the CDsampling package. Additionally, W_func_GLM() function can be used to find the diagonal elements of the matrix $\mathbf{W}$ in the Fisher information matrix (1) of GLM.
@@
We revisit the trauma_data example previously analyzed in Section 3.2. Consider a research study under a cumulative logit npo model with $J=5$ response levels and covariates $(x_{i1},x_{i2})\in\{(1,0),(2,0),(3,0),(4,0),(1,1),(2,1),(3,1),$ (4,1)}(4,1)\}. The model can be written as the following four equations:
@@
where $i=1,\dots,8$. We assume that the true parameters $\boldsymbol{\beta}=(\hat{\beta}_{11},\hat{\beta}_{12},\hat{\beta}_{13},\hat{\beta}_{21},\hat{\beta}_{22},\hat{\beta}_{23},\hat{\beta}_{31},$ $\hat{\beta}_{32},\hat{\beta}_{33},\hat{\beta}_{41},\hat{\beta}_{42},\hat{\beta}_{43})^{\top}$ == (4.047,0.131,4.214,2.225,0.376,3.519,0.302,0.237,2.420,1.386,(-4.047,-0.131,4.214,-2.225,-0.376,3.519,-0.302,-0.237,2.420,1.386, $0.120,1.284)^{\top}$, and the approximate allocation for the eight design points is $\mathbf{w}=(1/8,\ldots,1/8)^{\top}$.
@@
> J=5; p=12; m=8;> beta = c(-4.047, -0.131, 4.214, -2.225, -0.376, 3.519, -0.302, -0.237,  2.420, 1.386,-0.120,  1.284)> Xi=rep(0,J*p*m); dim(Xi)=c(J,p,m)> Xi[,,1] = rbind(c( 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 1, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 1, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,2] = rbind(c( 1, 2, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 2, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 2, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 2, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,3] = rbind(c( 1, 3, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 3, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 3, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 3, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,4] = rbind(c( 1, 4, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 4, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 4, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 4, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,5] = rbind(c( 1, 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 1, 1, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 1, 1, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 1),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,6] = rbind(c( 1, 2, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 2, 1, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 2, 1, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 2, 1),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,7] = rbind(c( 1, 3, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 3, 1, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 3, 1, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 3, 1),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,8] = rbind(c( 1, 4, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 4, 1, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 4, 1, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 4, 1),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> alloc = rep(1/8,m)> F_func_MLM(w=rep(1/8,m), beta=beta, X=Xi, link=’cumulative’)Dimensions: 12 x 12Matrix:------------------------------------------------------------      [,1]        [,2]        [,3]        [,4]        [,5]        [,6] [1,]  0.44505694  1.37915564  0.43609135 -0.37247296 -1.21252053 -0.36379349 [2,]  1.37915564  4.78410934  1.35694145 -1.21252053 -4.31436344 -1.19085831 [3,]  0.43609135  1.35694145  0.43609135 -0.36379349 -1.19085831 -0.36379349 [4,] -0.37247296 -1.21252053 -0.36379349  0.51192600  1.55177413  0.48018678 [5,] -1.21252053 -4.31436344 -1.19085831  1.55177413  5.31193908  1.48241981 [6,] -0.36379349 -1.19085831 -0.36379349  0.48018678  1.48241981  0.48018678 [7,]  0.00000000  0.00000000  0.00000000 -0.09154268 -0.22027625 -0.07471168 [8,]  0.00000000  0.00000000  0.00000000 -0.22027625 -0.64323991 -0.18445802 [9,]  0.00000000  0.00000000  0.00000000 -0.07471168 -0.18445802 -0.07471168[10,]  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000[11,]  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000[12,]  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000      [,7]        [,8]        [,9]        [,10]       [,11]       [,12] [1,]  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000 [2,]  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000 [3,]  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000  0.00000000 [4,] -0.09154268 -0.22027625 -0.07471168  0.00000000  0.00000000  0.00000000 [5,] -0.22027625 -0.64323991 -0.18445802  0.00000000  0.00000000  0.00000000 [6,] -0.07471168 -0.18445802 -0.07471168  0.00000000  0.00000000  0.00000000 [7,]  0.29320484  0.71978889  0.16205254 -0.10435894 -0.25399393 -0.06194131 [8,]  0.71978889  2.13312603  0.41294396 -0.25399393 -0.74842743 -0.15305771 [9,]  0.16205254  0.41294396  0.16205254 -0.06194131 -0.15305771 -0.06194131[10,] -0.10435894 -0.25399393 -0.06194131  0.17861575  0.45287619  0.06925715[11,] -0.25399393 -0.74842743 -0.15305771  0.45287619  1.37180187  0.17395849[12,] -0.06194131 -0.15305771 -0.06194131  0.06925715  0.17395849  0.06925715------------------------------------------------------------
@@
This suggests that in the feasible allocation set $S_{0}=\{(w_{1},w_{2},w_{3})^{\top}$ $\in$ $\mathbb{R}^{3}\mid w_{i}\geq 0,i=1,\ldots,3;\sum_{i=1}^{3}w_{i}=1\}$, the D-optimal allocation is $\mathbf{w}=(w_{1},w_{2},w_{3})=(1/3,1/3,1/3)$. $\Box$
@@
We also need to find the upper boundary and lower boundary functions for $r_{i1},r_{i2}]$ in step $3^{\circ}$ of the constrained lift-one algorithm to meet the conditions in $S$. The $r_{i1}$ and $r_{i2}$ can be obtained as follows:
@@
To determine the EW D-optimal allocation with the independent uniform priors $\beta_{0}\sim$ uniform(2,2)(-2,2), $\beta_{1}\sim$ uniform(1,5)(-1,5), $\beta_{21}\sim$ uniform(1,5)(-1,5), and $\beta_{22}\sim$ uniform(-1,5), as outlined in Section 3.1, we utilize the same codes provided in Section 3.1:
> library(cubature)> unif.prior <- rbind(c(-2, -1, -1, -1), c(2,  5,  5, 5))> W.EW.unif = matrix(rep(0,6))> for (i in 1:6){+   x = matrix((cbind(1, unique(X)))[i,])+   W.EW.unif[i] = hcubature(function(beta) dunif(beta[1], min=unif.prior[1,1],max=unif.prior[2,1])*dunif(beta[2], min=unif.prior[1,2],max=unif.prior[2,2])*dunif(beta[3], min=unif.prior[1,3],max=unif.prior[2,3])*dunif(beta[4], min=unif.prior[1,4], max=unif.prior[2,4])*(exp(x[1]*beta[1]+x[2]*beta[2]+x[3]*beta[3]+x[4]*beta[4])/(1+exp(x[1]*beta[1]+x[2]*beta[2]+x[3]*beta[3]+x[4]*beta[4]))ˆ2), lowerLimit = unif.prior[1,],upperLimit  = unif.prior[2,])\$integral+ }> set.seed(602)> approximate_design_EW = liftone_constrained_GLM(X=X, W=W.EW.unif,  g.con=g.con,+ g.dir=g.dir, g.rhs=g.rhs, lower.bound=lower.bound, upper.bound=upper.bound,+ reltol=1e-12, maxit=100, random=TRUE, nram=12, w00=NULL, epsilon=1e-12)> exact_design_EW = approxtoexact_constrained_func(n=200, w=approximate_design_EW\$w,+ m=6, beta=beta, link=’logit’, X=X, Fdet_func=Fdet_func_GLM, iset_func=iset_func_trial)
@@
> nsimu=100> p=3 # 3 covariates (excluding intercept)> set.seed(666)> seeds <- sample(1:10000, nsimu)   #random seeds> # matrix for recording beta estimates> beta.estimate <- rep(NA, nsimu*(p+1)*(5))> dim(beta.estimate) = c(nsimu, p+1, 5) # [1] 100   4   8> dimnames(beta.estimate)[[2]] = paste0("beta",0:p,seq="")> dimnames(beta.estimate)[[3]] <- c(’full’, "SRSWOR", "Unif", "local_Dopt", "EW_Unif")> # generate X for original 500 patients> # gender group (0 for female and 1 for male)> gender = c(rep(0,50),rep(0,40),rep(0,10),rep(1,200),rep(1,150),rep(1,50))> # age group (0 for 18˜25, 1 for 26˜64, and 2 for 65 or above)> age1 = c(rep(0,50), rep(1,40), rep(0,10), rep(0,200), rep(1,150), rep(0,50))> age2 = c(rep(0,50), rep(0,40), rep(1,10), rep(0,200), rep(0,150), rep(1,50))> X.original = cbind(gender, age1, age2)> #probability of Y = 1 given X using coefficient beta under logistic model> prob = exp(beta %*% t(cbind(1,X.original))) / (1+exp(beta %*% t(cbind(1,X.original))))> label = rep(0, length(X.original[,1])) #create label to store category of the population data> for(k in 1:length(X.original[,1])){+   for (m in 1:m){+     if((X.original[k,]==unique(X.original)[m,])[1] & (X.original[k,]==unique(X.original)[m,])[2]& (X.original[k,]==unique(X.original)[m,])[3]){+       label[k] = m+     }+   }+ }> n = tabulate(label) #number of subjects available in each category> sampling_func = function(label, n.sample, X, s, n, m){+   J = 0+   j = rep(0, m)+   c = rep(0, m)+   isample = vector()+   while(J < n.sample){+     for(k in 1:length(X[,1])){+       i = label[k]+       c[i] = c[i] + 1+       if(j[i]<s[i]){+         if(runif(1) < (s[i] - j[i])/(n[i] - c[i] + 1)){+           j[i]=j[i]+1+           J = J + 1+           isample = c(isample, k)+         }+       }+     }+   }+   return(isample)+ }> s_unif = unif_design\$allocation #num to sample in constrained uniform from each category> s_localD = exact_design\$allocation #num to sample in local D-optimal from each category> s_EWD = exact_design_EW\$allocation #num to sample in EW D-optimal from each category> for(isimu in 1:nsimu){+   set.seed(seeds[isimu])+   ##Generate Y response under logistic model+   Y = rbinom(n=length(X.original[,1]), size=1, prob = prob)+   #Estimate beta with full data+   beta.estimate[isimu,,"full"] <- glm(Y˜X.original,family = "binomial")\$coefficients+   #SRSWOR sample+   isample1 <- sample(x=1:500, size=200);+   X1=X.original[isample1,];+   Y1=Y[isample1];+   #Estimate beta with SRSWOR sampling+   beta.estimate[isimu,,"SRSWOR"] <- glm(Y1˜X1,family = "binomial")\$coefficients+   #Uniform sample+   #stratified function returns s[k] (num of samples plan to collect for ith category)+   isample2 = sampling_func(label=label, n.sample=200, X=X.original, s=s_unif, n=n, m=m)+   X2=X.original[isample2,];+   Y2=Y[isample2];+   #Estimate beta with Unif sampling+   beta.estimate[isimu,,"Unif"] <- glm(Y2˜X2,family = "binomial")\$coefficients+   #Local D-opt Sample+   isample3 = sampling_func(label=label, n.sample=200, X=X.original, s=s_localD, n=n,m=m)+   X3=X.original[isample3,];+   Y3=Y[isample3];+   #Estimate beta with local D-optimal sampling+   beta.estimate[isimu,,"local_Dopt"] <- glm(Y3˜X3,family = "binomial")\$coefficients+   #EW D-opt Sample+   isample4 = sampling_func(label=label, n.sample=200, X=X.original, s=s_EWD, n=n, m=m)+   X4=X.original[isample4,];+   Y4=Y[isample4];+   #Estimate beta with EW D-optimal sampling+   beta.estimate[isimu,,"EW_Unif"] <- glm(Y4˜X4,family = "binomial")\$coefficients+ }